Circular No. 8990

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2008 TC_{3}

A small asteroidal object (absolute magnitude H = 30.4, suggesting a few meters in size) discovered by the Mt. Lemmon Survey (observer R. A. Kowalski) on Oct. 6.28 UT at Δ = 0.0033 AU (∼ 1.27 the moon’s distance from the earth) will enter the earth’s atmosphere over northern Sudan (according to S. Chesley, Jet Propulsion Laboratory) around Oct. 7.115, moving west to east. Prior to entering the atmosphere, it may be around visual mag 11. Astrometry, orbital elements, and ephemerides are given on MPECs 2008-T50, 2008-T51, and 2008-T52.

QY MUSCAE W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag approximately 8.6) on a pair of Technical Pan photographs taken on Sept. 28.998 UT with a 85-mm camera lens (+ orange filter). The new object is located at α = 13^{h}16^{m}30^{s}, δ = −67^{o}37′.0 (equinox 2000.0); additional magnitudes from Liller: Sept. 15.023, [11.5; Oct. 4.032, 8.9. Liller adds that a weak spectrogram taken on Oct. 4.015 with a Schmidt camera of focal length 300 mm (+ orange filter + Technical Pan film + 100-mm × 100-mm transmission grating with 75 grooves/mm) shows a single broad emission line at the expected position of Hα (width at least ≈ 230 nm). V. Tabur, Wanniassa, Australian Capital Territory, reports that his unfiltered CCD images, taken with an intentionally significantly defocused Nikon ED 77-mm-aperture f/2.8 camera lens (+ SBIG ST8XE CCD camera) yields the following position (estimated uncertainty ± 4′′) for the apparent nova: α = 13^{h}16^{m}36^{s}.22, δ = −67^{o}36′50′′.7; while the variable is involved with a nearby star due to the defocused images, he provides the following magnitudes for the apparent nova: Sept. 19.418, [11.6; 21.378, 9.9; 23.401, 9.8; 24.378, 9.7; 25.378, 9.2; 27.389, 8.7; 28.383, 8.6; 29.436, 8.5 (cloud interfering); 30.397, 8.1; Oct. 1.395, 8.3. P. Williams, Heathcote, N.S.W., reports visual mag 8.9 on Oct. 2.440. M. Templeton, AAVSO, forwards the following position end figures and magnitudes from P. Nelson, Ellinbank, Victoria, Australia: 36^{s}.44, 47′′.8; B = 10.26, V = 9.31, at Oct. 2.44. B. A. Skiff, Lowell Observatory, reports that a USNO-B1.0-catalogue star of blue mag 19.9 and red mag 17.5 has position end figures 36^{s}.47, 47′′.9, noting that a star is also present with 0′′.5 of Nelson’s position on a plate taken via the U.K. Schmidt Telescope Hα survey. E. Kazarovets reports that the GCVS team assigns the designation QY Mus to this object.

2008 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green